OpenStudy (anonymous):

why is 1^1/2 equal to 1?

OpenStudy (anonymous):

Power half and square root are the same (power third would be cube root)

OpenStudy (luffingsails):

By definition 1*1 = 1, therefore 1^2=1. So, 1^(1/2) = 1.

OpenStudy (anonymous):

Like estudier said 1^(1/2) = 1 It is the same thing as writing √1 = 1